OpenStudy (anonymous):

HELP! The ratio of areas between two similar triangles is 1:4. if one side of the smaller triangle is 2 units, find the measure of the corresponding side of the other triangle.

OpenStudy (raden):

(side 1/side 2)^2 = area 1/area 2 let the measure of the bigger side is x, so (2/x)^2 = 1/4 4/x^2 = 1/4 cross multiply, get x^2 = 16 x = sqrt(16) x = 4

OpenStudy (anonymous):

thanks Raden! thanks a lot!!

OpenStudy (raden):

welcome